OpenStudy (anonymous):

Which has the greatest value? Answer 3 to the fourth power 5 to the third power ten to the second power seventeen to the zero power

hartnn (hartnn):

no its not

hartnn (hartnn):

\[3^4=81\] \[5^3=125\] \[10^2=100\] \[17^0=1\] now can u tell which one is greatest??